# Deep extragalactic H i survey of the COSMOS field with FAST

### v1

@@ Abstract
We present a deep H i survey at L-band conducted with the Five-hundred-meter Aperture Spherical radio Telescope (FAST) over the COSMOS field. This survey is strategically designed to overlap with the MIGHTEE COSMOS field, aiming to combine the sensitivity of the FAST and high-resolution of the MeerKAT. We observed the field with FAST for 11 hours covering square degrees, and reduced the raw data to H i spectral cubes over the frequency range 1310-1420 MHz. The FAST-H i data reach a median 3 column density of cm-2 over a channel width, allowing for studies of the distribution of H i gas in various environments, such as in galaxies, the Circum-Galactic Medium (CGM) and Intergalactic Medium (IGM). We visually searched the spectral cubes for H i sources, and found a total of 80 H i detections, of which 56 have been cross-matched with the MIGHTEE-H i catalogue. With the cross-matched sources, we compare their H i masses and find that the total H i mass fraction in the IGM and CGM surrounding the galaxy pairs is statistically higher than the H i fraction surrounding the isolated galaxies by a difference of 13%, indicating that the CGM and IGM associated with interacting systems are richer in neutral hydrogen compared to those around isolated galaxies in the local Universe. We also describe several FAST-MeerKAT synergy projects, highlighting the full potential of exploiting both single-dish and interferometric observations to study the distribution and evolution of the diffuse H i gas.
@@
With the sensitivity of the Five-hundred-meter Aperture Spherical Radio Telescope (FAST; Nan et al., 2011), we now have a facility with the potential to detect these faint H i emissions within a reasonable integration time. However, due to the large beam size (3 arcmin) of FAST, disentangling the IGM/CGM from the H i gas within individual galaxies is limited to the very local nearby Universe. Fortunately, with the excellent angular resolution (10 arcsec) of the MeerKAT radio telescope (Jonas & MeerKAT Team, 2016), we can identify the H i within these galaxies directly, and then use a combination of data from FAST and MeerKAT to extract the IGM/CGM signal at the cm-2 regime thoroughly. To achieve this, we carried out a series of FAST observations in L-band across the COSMOS field, overlapping with one of the MeerKAT International GHz Tiered Extragalactic Exploration (MIGHTEE; Jarvis et al., 2016) fields.
@@ 2.2 Ancillary data
However, we use the latest MIGHTEE Data Release 1 which is observed with the 32k channels (Heywood et al. submitted) for comparing the H i correlator mode. We smooth a cubelet surrounding each source in the Early Science catalogue provided by Ponomareva et al. (2023), and clip it at a 3 threshold as a mask for removing the noise following (Ponomareva et al., 2021). The MeerKAT 32k correlator mode has 32,768 channels with a spectral resolution of 26.1 kHz (i.e. 5.5 kms-1 at 1420 MHz). The COSMOS field was observed in this mode for a total of 158h tracks in a tightly-dithered mosaic that spans 2 deg2, and each pointing is imaged with robustness parameters of 0.0 and 0.5 (Briggs, 1995) following up on two rounds of RFI flagging, self-calibration, and the visibility-domain continuum subtraction. Then all pointings are brought together for homogenisation, mosaicking, and a further image-domain continuum subtraction (Heywood et al. submitted and Jarvis et al., 2024). We use the data product with the robustness of 0.5 (i.e. angular resolution 16 arcseconds across 1300-1420 MHz) to have a well balanced combination of the image resolution and sensitivity for detecting the H i gas within galaxies. We note that a robust catalogue is currently not available for the 32k data, which is why we use the Early Science catalogue to match to the galaxy positions.
@@ 3.2.1 Cross matching
After cross-matching the sample of 80 FAST-H i detections with the MIGHTEE-H i Early Science catalogue (Ponomareva et al., 2023) (but using the measurements from Date Release 1; Heywood et al. submitted) by setting the maximum angular separation of 3 arcminites and velocity offset of 200 , we find 56 cross-matched sources. A comparison of their spatial distribution is shown in Figure 3. The rate of finding a MIGHTEE-H i counterpart is clearly higher in the centre region with the highest sensitivity. We plot their H i masses against redshift in the left panel of Figure 4, and the H i flux against in the right panel to investigate the differences between the FAST-H i and MIGHTEE-H i detections. It shows that the non MIGHTEE-detected sources are dominated by the low-flux or narrow-line width H i galaxies as would be expected given the coarse spectrum resolution of 44 for the MIGHTEE-H i Early Science data and the limited sensitivity. However, there are a few non cross-matched H i sources that have a moderate amount of H i gas and a velocity width larger than 150 in the FAST data. In other words, some non cross-matched H i sources appear to be relatively bright galaxies, and are detected in FAST-H i but not in the MIGHTEE-H i 4k catalogue. This difference is likely indicating the fundamental difference between MeerKAT and FAST as an interferometer is generally less sensitive to extended/diffuse emission while a single dish telescope is sensitive to both compact and extended emission. The source finding with the MIGHTEE-H i 32k data will help further examine these sources. There are also some sources that are detected by MeerKAT but not FAST from our initial visual source finding. The presence of systematics in the FAST data complicates baseline subtraction, potentially contributing to the “missing sources” from the FAST data, especially for the faint ones. By using the positions from the MIGHTEE-H i detections, we extract the FAST-H i fluxes for the missing ones and find that these non-detections are indeed very faint with signal to noise ratios less than . Their spectra and masses are shown in Appendix B.

### v2

@@ Abstract
We present a deep H i survey at L-band conducted with the Five-hundred-meter Aperture Spherical radio Telescope (FAST) over the COSMOS field. This survey is strategically designed to overlap with the MIGHTEE COSMOS field, aiming to combine the sensitivity of the FAST and high-resolution of the MeerKAT. We observed the field with FAST for approximately 11 hours covering square degrees, and reduced the raw data to H i spectral cubes over the frequency range 1310-1420 MHz. The FAST-H i data reach a median 3 column density of cm-2 over a channel width, allowing for studies of the distribution of H i gas in various environments, such as in galaxies, the Circum-Galactic Medium (CGM) and Intergalactic Medium (IGM). We visually searched the spectral cubes for H i sources, and found a total of 80 H i detections, of which 56 have been cross-matched with the MIGHTEE-H i catalogue. With the cross-matched sources, we compare their H i masses and find that the total H i mass fraction in the IGM and CGM surrounding the galaxy pairs is statistically higher than the H i fraction surrounding the isolated galaxies by a difference of 13%, indicating that the CGM and IGM associated with interacting systems are richer in neutral hydrogen compared to those around isolated galaxies in the local Universe. We also describe several FAST-MeerKAT synergy projects, highlighting the full potential of exploiting both single-dish and interferometric observations to study the distribution and evolution of the diffuse H i gas.
@@
With the sensitivity of the Five-hundred-meter Aperture Spherical Radio Telescope (FAST; Nan et al., 2011), we now have a facility with the potential to detect these faint H i emissions within a reasonable integration time, as indicated, for example, by the FAST All Sky HI survey (FASHI; Zhang et al., 2024). However, due to the large beam size (3 arcmin) of FAST, disentangling the IGM/CGM from the H i gas within individual galaxies is limited to the very local nearby Universe. Fortunately, with the excellent angular resolution (10 arcsec) of the MeerKAT radio telescope (Jonas & MeerKAT Team, 2016), we can identify the H i within these galaxies directly, and then use a combination of data from FAST and MeerKAT to extract the IGM/CGM signal at the cm-2 regime thoroughly. To achieve this, we carried out a series of FAST observations in L-band across the COSMOS field, overlapping with one of the MeerKAT International GHz Tiered Extragalactic Exploration (MIGHTEE; Jarvis et al., 2016) fields.
@@ 2.2 Ancillary data
However, we use the latest MIGHTEE Data Release 1 which is observed with the 32k channels (Heywood et al., 2024) for comparing the H i correlator mode. We smooth a cubelet surrounding each source in the Early Science catalogue provided by Ponomareva et al. (2023), and clip it at a 3 threshold as a mask for removing the noise following (Ponomareva et al., 2021). The MeerKAT 32k correlator mode has 32,768 channels with a spectral resolution of 26.1 kHz (i.e. 5.5 kms-1 at 1420 MHz). The COSMOS field was observed in this mode for a total of 158h tracks in a tightly-dithered mosaic that spans 2 deg2, and each pointing is imaged with robustness parameters of 0.0 and 0.5 (Briggs, 1995) following up on two rounds of RFI flagging, self-calibration, and the visibility-domain continuum subtraction. Then all pointings are brought together for homogenisation, mosaicking, and a further image-domain continuum subtraction (Heywood et al., 2024; Jarvis et al., 2024). We use the data product with the robustness of 0.5 (i.e. angular resolution 16 arcseconds across 1300-1420 MHz) to have a well balanced combination of the image resolution and sensitivity for detecting the H i gas within galaxies. We note that a robust catalogue is currently not available for the 32k data, which is why we use the Early Science catalogue to match to the galaxy positions.
@@ 3.2.1 Cross matching
After cross-matching the sample of 80 FAST-H i detections with the MIGHTEE-H i Early Science catalogue (Ponomareva et al., 2023), but using the measurements from Date Release 1 (Heywood et al., 2024), by setting the maximum angular separation of 3 arcminites and velocity offset of 200 , we find 56 cross-matched sources. A comparison of their spatial distribution is shown in Figure 3. The rate of finding a MIGHTEE-H i counterpart is clearly higher in the centre region with the highest sensitivity. We plot their H i masses against redshift in the left panel of Figure 4, and the H i flux against in the right panel to investigate the differences between the FAST-H i and MIGHTEE-H i detections. It shows that the non MIGHTEE-detected sources are dominated by the low-flux or narrow-line width H i galaxies as would be expected given the coarse spectrum resolution of 44 for the MIGHTEE-H i Early Science data and the limited sensitivity. However, there are a few non cross-matched H i sources that have a moderate amount of H i gas and a velocity width larger than 150 in the FAST data. In other words, some non cross-matched H i sources appear to be relatively bright galaxies, and are detected in FAST-H i but not in the MIGHTEE-H i 4k catalogue. This difference is likely indicating the fundamental difference between MeerKAT and FAST as an interferometer is generally less sensitive to extended/diffuse emission while a single dish telescope is sensitive to both compact and extended emission. The source finding with the MIGHTEE-H i 32k data will help further examine these sources. There are also some sources that are detected by MeerKAT but not FAST from our initial visual source finding. The presence of systematics in the FAST data complicates baseline subtraction, potentially contributing to the “missing sources” from the FAST data, especially for the faint ones. By using the positions from the MIGHTEE-H i detections, we extract the FAST-H i fluxes for the missing ones and find that these non-detections are indeed very faint with signal to noise ratios less than . Their spectra and masses are shown in Appendix B.